how much time will it take for a ball to reach its maximum height it it was given initial speed of 120ft/s and its angle of inclination is 32 degrees?


1
Expert's answer
2021-07-14T10:16:09-0400

Let the initial speed be "v_0 = 120ft/s \approx 36.58m/s", and the angle of inclination "\theta = 32\degree". Then the time it will take for a ball to reach its maximum height is (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"t = \dfrac{v_0\sin\theta}{g}\\
t = \dfrac{36.58\cdot \sin32\degree}{9.81}\approx 1.98s"

Answer. 1.98s.